\begin{thebibliography}{99}

\bibitem{Addison} A. W. Addison, A series representation for Euler's constant,
Amer. Math. Monthly {\bf 74} (1967) 823--824.

\bibitem{AS} J.-P. Allouche and J. Shallit,
Sums of digits and the Hurwitz zeta function,
in Analytic number theory (Tokyo, 1988), Lecture Notes in Math. {\bf 1434}, Springer, 
Berlin, 1990, pp. 19--30. 

\bibitem{ASreg} J.-P. Allouche and J. Shallit, The ring of $k$-regular sequences,
Theoret. Comput. Sci. {\bf 98} (1992) 163--197.

\bibitem{ASreg2} J.-P. Allouche and J. Shallit, The ring of $k$-regular
sequences II, Theoret. Comput. Sci. {\bf 307} (2003) 3--29. 

\bibitem{Behrmann} A. Behrmann, Problem 5460, Amer. Math. Monthly {\bf 74}
(1967) 206.

\bibitem{BerBow} B. C. Berndt and D. C. Bowman, Ramanujan's short unpublished 
manuscript on integrals and series related to Euler's constant, in 
Constructive, Experimental and Nonlinear Analysis, M. Th\'era, ed., 
American Mathematical Society, Providence, RI, 2000, pp. 19--27.

\bibitem{Catalan} E. Catalan, Sur la constante d'Euler et la fonction de Binet,
J. Math. Pures Appl. {\bf 1} (1875) 209--240.

\bibitem{ErMaObTr} A. Erd\'elyi, W. Magnus, F. Oberhettinger, and F. G. Tricomi,
Higher transcendental functions, Bateman manuscript project,
Vol.~1, McGraw-Hill, New York, 1953.

\bibitem{KAH} L. F. Klosinski, G. L. Alexanderson, and A. P. Hillman, The William
Lowell Putnam mathematical competition, Amer. Math. Monthly {\bf 89} (1982) 
679--686.

\bibitem{Knuth} D. E. Knuth, Fundamental algorithms, the art of computer
programming, Vol.~1, Addison-Wesley, Reading, 1981.

\bibitem{Lint} J. H. van Lint, Solution to Problem 5460, Amer. Math. Monthly 
{\bf 75} (1968) 202.

\bibitem{MM} P. Morton and W. Mourant, Paper folding, digit patterns, and groups of 
arithmetic fractals, Proc. London Math. Soc. {\bf 59} (1989) 253--293.

\bibitem{Shallit} J. Shallit, Advanced Problem 6450, Amer. Math. Monthly {\bf 91}
(1984) 59--60. Solution appeared in Amer. Math. Monthly {\bf 92} (1985) 513--514.

\bibitem{Sondow1} J. Sondow, Double integrals for Euler's constant and
$\ln(4/\pi)$ and an analog of Hadjicostas's formula, Amer. Math. Monthly 
{\bf 112} (2005) 61--65.

\bibitem{Sondow2} J. Sondow, New Vacca-type rational series for Euler's constant
and its ``alternating'' analog $\ln \frac{4}{\pi}$, preprint 2005, available at

\noindent
{\tt http://arxiv.org/abs/math.NT/0508042}

\bibitem{SondZudi} J. Sondow and W. Zudilin, Euler's constant, $q$-logarithms, 
and formulas of Ramanujan and Gosper, Ramanujan J., to appear. 

\bibitem{Vacca} G. Vacca, A new series for the Eulerian constant 
$\gamma = \cdot577\ldots$, Quart. J. Math. {\bf 41} (1910) 363--364.

\bibitem{WW} E. T. Whittaker and G. N. Watson, A course of modern analysis, 
Cambridge University Press, Cambridge, 1978.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{theorem,amssymb}

\advance \topmargin by -\headheight
\advance \topmargin by -\headsep
\evensidemargin \oddsidemargin
\author{J.-P. Allouche\thanks{partially supported 
by MENESR, ACI NIM 154 Num\'eration.} \\
CNRS, LRI, B\^atiment 490 \\
F-91405 Orsay Cedex \\
France\\
{\tt allouche@lri.fr} \\
\and
J. Shallit \\
School of Computer Science, University of Waterloo \\
Waterloo, Ontario  N2L 3G1 \\
Canada \\
{\tt shallit@graceland.uwaterloo.ca} \\
\and
J. Sondow \\
209 West 97th Street \\
New York, NY 10025  \\
USA\\
{\tt jsondow@alumni.princeton.edu} \\
}

\title{Summation of Series Defined by Counting Blocks of Digits}

\date{ }

\def \proof{\bigbreak\noindent{\it Proof.\ \ }}

\def \endpf{{\ \ $\Box$ \medbreak}}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\theorembodyfont{\rm}

\newtheorem{remark}{Remark}
\newtheorem{example}{Example}

\begin{document}

\maketitle

\begin{abstract}
We discuss the summation of certain series defined by counting blocks of 
digits in the $B$-ary expansion of an integer. For example, if $s_2(n)$ 
denotes the sum of the base-$2$ digits of $n$, we show that 
$\sum_{n \geq 1} s_2(n)/(2n(2n+1)) = (\gamma + \log \frac{4}{\pi})/2$.
We recover this previous result of Sondow and provide several generalizations.

\medskip

\noindent
{\it MSC}:  11A63, 11Y60. 
\end{abstract}

\section{Introduction} 

A classical series with rational terms, known as Vacca's series
\cite{Vacca} or in an equivalent integral form as Catalan's integral
\cite{Catalan} (see also \cite{BerBow} and \cite{SondZudi}), evaluates 
to Euler's constant $\gamma$:
$$
\gamma = \sum_{n \geq 1} \frac{(-1)^n}{n} 
\left\lfloor \frac{\log n}{\log 2} \right\rfloor = 
\int_0^1 \frac{1}{1+x} \sum_{n \geq 1} x^{2^n -1}dx.
$$
In a recent paper \cite{Sondow2} Sondow gave the following two 
formulas:
$$
\gamma^{\pm} = \sum_{n \geq 1} \frac{N_1(n) \pm N_0(n)}{2n(2n+1)}
$$
where $\gamma^+ = \gamma$ is the Euler constant, 
$\gamma^- = \log \frac{4}{\pi}$ is the ``alternating Euler 
constant'' \cite{Sondow1}, and $N_1(n)$ (resp. $N_0(n)$) is 
the number of $1$'s (resp. $0$'s) in the binary expansion of the 
integer $n$. The series for $\gamma^+ = \gamma$ is equivalent to 
Vacca's. The formulas for $\gamma^{\pm}$ show in particular that
$$
\sum_{n \geq 1} \frac{s_2(n)}{2n(2n+1)} =
\frac{\gamma + \log \frac{4}{\pi}}{2}
$$
where $s_2(n)$ is the sum of the binary digits of the integer $n$.

This last formula reminds us of one of the problems posed at the 1981 Putnam 
competition \cite{KAH}: Determine whether or not 
$$
\exp\left(\sum_{n \geq 1} \frac{s_2(n)}{n(n+1)}\right)
$$
is a rational number. 
In fact, $\sum \frac{s_2(n)}{n(n+1)} = 2\log 2$. A generalization 
was proven by Shallit \cite{Shallit}, where the base $2$ is replaced by
any integer base $B \geq 2$. A more general result, where the sum of digits is
replaced by the function
$N_{w,B}(n)$, which counts the number of occurrences of the
block $w$ in the $B$-ary expansion of the integer $n$, was given 
by Allouche and Shallit \cite{AS}.

The purpose of the present paper is to show that the result of \cite{Sondow2} 
cited above can be deduced from a general lemma in \cite{AS}. Furthermore, we 
sum the series
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)}
\ \ \ \mbox{\rm and} \ \ \
\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)(2n+2)},
$$
thus generalizing Corollary~1 in \cite{Sondow2} and a series for Euler's 
constant in \cite{Behrmann, Addison, Lint} (dated February 1967, 
August 1967, February 1968), respectively. Finally, 
we indicate some generalizations of our results, including an extension 
to base $B > 2$, and a method for giving alternate proofs without using 
the general lemma from \cite{AS}.

\section{A general lemma}

The first lemma in this section is taken from \cite{AS}; for completeness we 
recall the proof. We also give two classical results presented as lemmas,
together with a new result (Lemma~\ref{mainlemma}).

We start with some definitions.
Let $B \geq 2$ be an integer. Let $w$ be a word on the alphabet of digits
$\{0, 1, \cdots, B-1\}$ (that is, $w$ is a finite block of digits).
We denote by $N_{w,B}(n)$ the number of (possibly overlapping) occurrences
of $w$ in the $B$-ary expansion of an integer $n > 0$, and we set 
$N_{w,B}(0) = 0$.

Given $w$ as above, we denote by $|w|$ the length of the word $w$ (i.e., 
if $w = d_1 d_2 \cdots d_k$, then $|w| = k$). Denote by $w^j$ the 
concatenation of $j$ copies of the word $w$.

Given $w$ and $B$ as above, we denote by $v_B(w)$ the value of $w$ when $w$ 
is interpreted as the base $B$-expansion (possibly with leading $0$'s) of an
integer.

\begin{remark}
The occurrences of a given word in the $B$-ary expansion of the integer $n$ 
may overlap. For example, $N_{11,2}(7) = 2$.

If the word $w$ begins with $0$, but $v_B(w) \neq 0$, then in computing
$N_{w,B}(n)$ we assume that the $B$-ary expansion of $n$ starts with an 
arbitrarily long prefix of $0$'s. If $v_B(w) = 0$ we use the usual $B$-ary 
expansion of $n$ without leading zeros.
For example, $N_{011,2}(3) = 1$
(write $3$ in base $2$ as $0\cdots011$) and $N_{0,2}(2) = 1$.
\end{remark}

\begin{lemma}[\cite{AS}]\label{AS}
Fix an integer $B \geq 2$, and let $w$ be a non-empty word on the alphabet
$\{0, 1, \cdots, B-1\}$. If $f: {\mathbb N} \to {\mathbb C}$ is a function
with the property that $\sum_{n \geq 1} |f(n)| \log n < \infty$, then
$$
\sum_{n \geq 1} N_{w,B}(n) \left(f(n) - \sum_{0 \leq j < B} f(Bn+j)\right)
= \sum f(B^{|w|} n + v_B(w)),
$$
where the last summation is over $n \geq 1$ if $w=0^j$ for some $j \geq 1$,
and over $n \geq 0$ otherwise.
\end{lemma}

\proof (See \cite{AS}.) 
As $N_{w,B}(n) \leq \lfloor\frac{\log n}{\log B}\rfloor + 1$, all series
$\sum N_{w,B}(un+v) f(un+v)$, where $u$ and $v$ are nonnegative integers,
are absolutely convergent. Let $\ell$ be the last digit of $w$, and
let $g :=B^{|w|-1}$. Then
$$
\sum_{n \geq 0} N_{w,B}(n) f(Bn+\ell) =
\sum_{0 \leq k < g} \sum_{n \geq 0} N_{w,B}(gn+k) f(Bgn+Bk+\ell)
$$
and
$$
\sum_{n \geq 0} N_{w,B}(Bn+\ell) f(Bn+\ell) =
\sum_{0 \leq k < g} \sum_{n \geq 0} N_{w,B}(Bgn+Bk+\ell) f(Bgn+Bk+\ell).
$$
Now, if either $n \neq 0$ or $v_B(w) \neq 0$, then for $k = 0,1,\ldots, g-1$
we have
$$
N_{w,B}(Bgn+Bk+\ell) - N_{w,B}(gn+k) =
\left\{
\begin{array}{ll}
1, \ &\mbox{\rm if } k = \lfloor \frac{v_B(w)}{B} \rfloor; \\
& \\
0, \ &\mbox{\rm otherwise}.
\end{array}
\right.
$$
On the other hand, if $n=0$ and $v_B(w)=0$ (hence $\ell=0$), then the 
difference equals $0$ for every $k \in \{0, 1, \cdots, g-1\}$. Hence
$$
\begin{array}{lll}
\displaystyle\sum_{n \geq 0} N_{w,B}(Bn+\ell) f(Bn+\ell) -
\displaystyle\sum_{n \geq 0} N_{w,B}(n) f(Bn+\ell) &=& 
\displaystyle\sum f\left(Bgn + B \lfloor\frac{v_B(w)}{B}\rfloor + \ell \right)
\\ 
&=& \displaystyle\sum f(B^{|w|} n + v_B(w)) \ \ \ \ (*),
\end{array}
$$
the last two summations being over $n \geq 0$ if $w$ is not of the form $0^j$, 
and over $n \geq 1$ if $w = 0^j$ for some $j \geq 1$. We then write
$$
\begin{array}{lll}
\displaystyle\sum_{n \geq 0} N_{w,B}(n) f(n) &=&
\displaystyle\sum_{0 \leq j < B} \ \sum_{n \geq 0} N_{w,B}(Bn+j) f(Bn+j) \\
&=& \displaystyle\sum_{j \in [0,B)\setminus\{\ell\}} 
\ \displaystyle\sum_{n \geq 0} N_{w,B}(Bn+j) f(Bn+j)
+ \displaystyle\sum_{n \geq 0} N_{w,B}(Bn+\ell) f(Bn+\ell)
\end{array}
$$
which together with $(*)$ gives
$$
\sum_{n \geq 0} N_{w,B}(n) \left(
f(n) - \sum_{0 \leq j < B} f(Bn+j) \right)
= \sum f(B^{|w|}n+v_B(w)).
$$
Since $N_{w,B}(0) = 0$, the proof is complete. \endpf

     Now let $\Gamma$ be the usual gamma function, let
$\Psi := \Gamma'/\Gamma$ be the logarithmic derivative of the gamma
function, let $\zeta(s)$ be the Riemann zeta function, let
$\zeta(s,x) := \sum_{n \geq 0} (n+x)^{-s}$ be the Hurwitz zeta function, and
let $\gamma$ denote Euler's constant.

\begin{lemma}\label{Hurwitz}
If $a$ and $b$ are positive real numbers, then
$$
\sum_{n \geq 1} \left(\frac{1}{an} - \frac{1}{an+b}\right) =
\frac{1}{b} + \frac{\gamma+\Psi(b/a)}{a} .
$$
\end{lemma}

\proof
We write
$$
\begin{array}{lll}
\displaystyle\sum_{n \geq 1} \left(\frac{1}{an} - \frac{1}{an+b}\right) &=& 
\displaystyle\lim_{s \to 1_+} \sum_{n \geq 1}
\left(\frac{1}{(an)^s} - \frac{1}{(an+b)^s}\right)
= \frac{1}{a} \displaystyle\lim_{s \to 1_+} \sum_{n \geq 1}
\left(\frac{1}{n^s} - \frac{1}{(n+\frac{b}{a})^s}\right) \\
&=&
\displaystyle\frac{1}{a} \displaystyle\lim_{s \to 1_+} \left(\zeta(s) - 
\zeta\left(s,\frac{b}{a}\right) + \left(\frac{a}{b}\right)^s\right) \\
&=& \displaystyle\frac{1}{b} +
\frac{1}{a} \displaystyle\lim_{s \to 1_+} 
\left(\left(\zeta(s) - \frac{1}{s-1}\right)
- \left(\zeta\left(s,\frac{b}{a}\right) - \frac{1}{s-1}\right)\right) \\
&=&
\displaystyle\frac{1}{b} + 
\frac{1}{a}\left(\gamma + \frac{\Gamma'(b/a)}{\Gamma(b/a)}\right)
= \frac{1}{b} + \frac{\gamma + \Psi(b/a)}{a}
\end{array}
$$
(see for example \cite[p. 271]{WW}). \endpf

\begin{lemma}\label{Gamma}
For $x > 0$ we have
$$
\sum_{r \geq 1} \left(\frac{x}{r} - \log \left(1+\frac{x}{r}\right)\right)
= \log x + \gamma x + \log \Gamma(x).
$$
\end{lemma}

\proof Take the logarithm of the Weierstra{\ss} product for $1/\Gamma(x)$
(see, for example, \cite[Section~1.1]{ErMaObTr} or 
\cite[Section~12.1]{WW}). \endpf

The next lemma in this section is the last step before proving 
our theorems.

\begin{lemma}\label{mainlemma}
Let $a$ and $b$ be positive real numbers. Then
$$
\sum_{n \geq 1} \left(\frac{1}{an} - \log \frac{an+1}{an}\right)
= \log \Gamma\left(\frac{1}{a}\right) + \frac{\gamma}{a} - \log a
$$
and
$$
\sum_{n \geq 0} \left(\frac{1}{an+b} - \log \frac{an+b+1}{an+b}\right)
= \log \Gamma\left(\frac{b+1}{a}\right) - \log \Gamma\left(\frac{b}{a}\right) 
- \frac{\Psi(b/a)}{a}.
$$
\end{lemma}

\proof The proof is straightforward. The first formula follows directly from 
Lemma~\ref{Gamma}. To prove the second, write the $n$th term of the series for
$n \geq 1$ as the following sum of $n$th terms of three absolutely 
convergent series:
$$
\frac{1}{an+b} - \frac{1}{an} - \frac{b}{an} 
+ \log\left(1+\frac{b}{an}\right) + \frac{b+1}{an}
- \log\left(1+\frac{b+1}{an}\right);
$$
then use Lemmas~\ref{Hurwitz} and \ref{Gamma}.
\endpf

\section{Two theorems}

In this section we give two theorems that are consequences of 
Lemma~\ref{AS}, and that generalize results in \cite{Sondow2} 
and \cite{Behrmann, Addison, Lint}. 

\begin{theorem}\label{degree2}
Let $w$ be a non-empty word on the alphabet $\{0, 1\}$, and let 
$\Psi$ denote the logarithmic derivative of the Gamma function.

\begin{itemize}
\item[(a)] If $v_2(w) = 0$, then
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)} =
\log \Gamma\left(\frac{1}{2^{|w|}}\right) + \frac{\gamma}{2^{|w|}}
- |w| \log 2.
$$

\item[(b)] If $v_2(w) \neq 0$, then
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)} =
\log \Gamma\left(\frac{v_2(w)+1}{2^{|w|}}\right) - 
\log \Gamma\left(\frac{v_2(w)}{2^{|w|}}\right) -
\frac{1}{2^{|w|}}\Psi\left(\frac{v_2(w)}{2^{|w|}}\right).
$$
\end{itemize}
\end{theorem}

\proof
Let 
$$
A_n := \frac{1}{n} - \log \frac{n+1}{n}
$$
for $n \geq 1$. 
Noting that $A_n - A_{2n} - A_{2n+1} = \frac{1}{2n(2n+1)}$, the theorem
follows from Lemma~\ref{AS} with $B=2$, and $f(n) := A_n$ for $n \geq 1$, 
together with Lemma~\ref{mainlemma}. \endpf

\begin{example}\label{gammaplusminus}

Taking $w = 0$ and $w = 1$, and recalling that 
$\Gamma(1/2) = \sqrt \pi$ and $\Psi(1/2) = - \gamma -2 \log 2$ 
by Gau{\ss}'s theorem (see for example \cite[p. 19]{ErMaObTr} or 
\cite[p. 94]{Knuth}), we get
$$
\sum_{n \geq 1} \frac{N_{0,2}(n)}{2n(2n+1)} =
\frac{1}{2} \log \pi + \frac{\gamma}{2} - \log 2
$$
and
$$
\sum_{n \geq 1} \frac{s_2(n)}{2n(2n+1)} =
\sum_{n \geq 1} \frac{N_{1,2}(n)}{2n(2n+1)} = 
- \frac{1}{2} \log \pi + \frac{\gamma}{2} + \log 2.
$$
These equalities imply the formulas in the Introduction:
$$
\sum_{n \geq 1} \frac{N_{1,2}(n) \pm N_{0,2}(n)}{2n(2n+1)} = \gamma^{\pm}
$$
where (following the notations of \cite{Sondow2}) $\gamma^+ := \gamma$
and $\gamma^- := \log \frac{4}{\pi}$, which is Corollary~1
of \cite{Sondow2}.

\end{example}

\begin{remark}
The formulas in Theorem~\ref{degree2} are analogous to those in 
\cite[p.~25]{AS}.  The analogy becomes more striking if one uses 
Gau{\ss}'s theorem to write all expressions of the form $\Psi(x)$, 
with $x$ a rational number in $(0,1]$, using only trigonometric 
functions, logarithms, and Euler's constant. 
\end{remark}

\begin{theorem}\label{degree3}
Let $w$ be a non-empty word on the alphabet $\{0, 1\}$.

\begin{itemize}
\item[(a)] If $v_2(w) = 0$, then
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)(2n+2)} =
\log \Gamma\left(\frac{1}{2^{|w|}}\right) + \frac{\gamma}{2^{|w|+1}}
- |w| \log 2 - \frac{1}{2^{|w|+1}}\Psi\left(\frac{1}{2^{|w|}}\right)
- \frac{1}{2}.
$$

\item[(b)] If $v_2(w) \neq 0$, then
$$
\begin{array}{ll}
\displaystyle\sum_{n \geq 1} \frac{N_{w,2}(n)}{2n(2n+1)(2n+2)} &=
\log \Gamma\displaystyle\left(\frac{v_2(w)+1}{2^{|w|}}\right) -
\log \Gamma\displaystyle\left(\frac{v_2(w)}{2^{|w|}}\right) \\
& \ \ \ \ \ - \displaystyle\frac{1}{2^{|w|+1}} 
\left(\Psi\left(\frac{v_2(w)}{2^{|w|}}\right)
+ \Psi\left(\frac{v_2(w)+1}{2^{|w|}}\right)\right).
\end{array}
$$
\end{itemize}
\end{theorem}

\proof 
Noting that $\frac{1}{2n(2n+1)} - \frac{1}{4} \cdot \frac{1}{n(n+1)} =
\frac{1}{2n(2n+1)(2n+2)}$, it suffices to use Theorem~\ref{degree2} 
and the following result, deduced from \cite[top of p. 26]{AS} in the 
case $B=2$.

{\it
\begin{itemize}
\item[(a)]
If $v_2(w) = 0$, then
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{n(n+1)} =
\frac{1}{2^{|w|-1}}
\left(
\Psi\left(\frac{1}{2^{|w|}}\right)
+ \gamma + 2^{|w|}
\right).
$$

\item[(b)] If $v_2(w) \neq 0$, then
$$
\sum_{n \geq 1} \frac{N_{w,2}(n)}{n(n+1)} =
\frac{1}{2^{|w|-1}}
\left(
\Psi\left(\frac{v_2(w)+1}{2^{|w|}}\right)
- \Psi\left(\frac{v_2(w)}{2^{|w|}}\right)
\right).
$$
\end{itemize}
}
\endpf

\begin{example}

Taking $w=0$ and $w=1$, we get
$$
\sum_{n \geq 1} \frac{N_{0,2}(n)}{2n(2n +1)(2n+2)} =
\frac{1}{2} \log \pi + \frac{\gamma}{2} - \frac{1}{2} \log 2 - \frac{1}{2}
$$
and
$$
\sum_{n \geq 1} \frac{s_2(n)}{2n(2n +1)(2n+2)} =
\sum_{n \geq 1} \frac{N_{1,2}(n)}{2n(2n +1)(2n+2)} = 
- \frac{1}{2} \log \pi + \frac{\gamma}{2} + \frac{1}{2} \log 2.
$$
Hence
$$
\sum_{n \geq 1} \frac{N_{1,2}(n) \pm N_{0,2}(n)}{2n(2n +1)(2n+2)} = \delta^{\pm}
$$
where $\delta^+ := \gamma -\frac{1}{2}$ and 
$\delta^- := \frac{1}{2} - \log \frac{\pi}{2}$, which are respectively 
a formula given in \cite{Behrmann, Addison, Lint} and a seemingly 
new companion formula.

\end{example}

\begin{remark}
As mentioned, all expressions of the form $\Psi(x)$, with
$x$ a rational number in $(0,1]$, can be written using only trigonometric
functions, logarithms, and Euler's constant.
\end{remark}

\section{Generalizations}\label{generalizations}

Several extensions or generalizations of our results are possible.
We give some of them in this section.

\subsection{Variation on $A_n$}

Instead of applying Lemma~\ref{AS} with $f(n) = A_n = \frac{1}{n} - \log \frac{n+1}{n}$ for $n \geq 1$,
we could replace $A_n$ with 
$$
A^{(k)}_n := \frac{1}{n+k} - \log \frac{n+1}{n}
$$
for $n \geq 1$, where $k$ is a nonnegative integer. Defining the (rational) 
function $Q^{(k)}$ by
$$
Q^{(k)}(n) := A^{(k)}_n - A^{(k)}_{2n} - A^{(k)}_{2n+1}
$$
and noting that summing $\sum_{n \geq 1} A^{(k)}_{an+b}$ boils down to 
summing $\sum_{n \geq 1} \left(\frac{1}{an+b+k} - \frac{1}{an+b}\right)$, 
which as in the proof of Lemma~\ref{Hurwitz} involves the Hurwitz zeta 
function, we obtain explicit formulas for the sum of the series 
$\sum_{n \geq 1} N_{w,2}(n) Q^{(k)}(n)$.

\subsection{Extension to base $B > 2$}\label{gal}

Lemma~\ref{AS} has been used above only for base $B=2$. There are
applications to other bases in \cite{AS}. We also note that the relation
among the $A_n$'s,
$$
A_n = \frac{1}{2n(2n+1)} + A_{2n} + A_{2n+1}
$$
for $n \geq 1$, can be generalized to base $B$. Namely,
$$
A_n = Q(n,B) + R(n,B)
$$
where
$$
Q(n,B) := \frac{1}{Bn(Bn+1)} + \frac{2}{Bn(Bn+2)} + 
\cdots + \frac{B-1}{Bn(Bn+B-1)}
$$
and
$$
R(n,B) := A_{Bn} + A_{Bn+1} + \cdots + A_{Bn+B-1}.
$$
This allows us to use Lemmas~\ref{AS} and \ref{mainlemma} to sum, for 
example, the series
$$
\sum_{n \geq 1} N_{w,3}(n) \frac{9n+4}{3n(3n+1)(3n+2)},
$$
since
$$
Q(n,3) = A_n - A_{3n} - A_{3n+1} - A_{3n+2} = \frac{9n+4}{3n(3n+1)(3n+2)}.
$$

\subsection{Weighted $A_n$'s}

In this section we consider a weighted form of the $A_n$'s.  
First we need to study a relation between sequences of real numbers.

\begin{lemma}\label{transform}
Let $(r_n)_{n \geq 1}$ and $(R_i)_{i \geq 1}$ be sequences of real numbers.
Set $r_0 := 0$ and $R_0 := 0$. Then the following two properties are equivalent:
\begin{itemize}
\item[(1)] for $i \geq 1$
$$
R_i = \sum_{k \geq 0} r_{\lfloor \frac{i}{2^k} \rfloor} =
r_i + r_{\lfloor\frac{i}{2}\rfloor} + r_{\lfloor\frac{i}{4}\rfloor} 
+ \cdots
$$
(note that this is actually a finite sum);

\item[(2)] for $n \geq 1$
$$
r_n = R_n - R_{\lfloor\frac{n}{2}\rfloor}.
$$
\end{itemize}
\end{lemma}

\proof 
The implication (1) $\Rightarrow$ (2) is easily seen by considering the 
cases $n$ even and $n$ odd. Likewise, for (2) $\Rightarrow$ (1) take 
$i$ even and $i$ odd.  \endpf

\begin{remark}
See \cite[Theorem~9]{ASreg2} for more about this relation.
\end{remark}

\begin{theorem}\label{weight}
Assume that $r_1, r_2, \ldots$ and $R_1, R_2, \ldots$ are real 
numbers related as in Lemma~\ref{transform}. Then the series 
$\sum |r_n| n^{-2}$ converges if and only if the series
$\sum |R_i| i^{-2}$ converges, and in this case we have
$$
S := \sum_{n \geq 1} r_n \left(\frac{1}{n} - \log \frac{n+1}{n}\right)
= \sum_{i \geq 1} \frac{R_i}{2i(2i+1)}
$$
\end{theorem}

\proof  First note that if the series $\sum |R_i|i^{-2}$ converges, 
then so does the series $\sum |r_n|n^{-2}$: use the expression for 
$r_n$ in terms of the $R_i$'s in Lemma~\ref{transform}. 
Now suppose that the series 
$\sum |r_n| n^{-2}$ converges. As before, let 
$A_n := \frac{1}{n} - \log \frac{n+1}{n}$. 
Then $0 < A_n < \frac{1}{n} - \frac{1}{n+1}$.
This implies that the series 
$S :=\sum r_n \left(\frac{1}{n} - \log \frac{n+1}{n}\right)$
is absolutely convergent. Now
$$
A_n = \frac{1}{2n(2n+1)} + A_{2n} + A_{2n+1}
$$ 
implies 
$$
A_n = \frac{1}{2n(2n+1)} + \frac{1}{4n(4n+1)} + \frac{1}{(4n+2)(4n+3)} 
+ A_{4n} + A_{4n+1} + A_{4n+2} + A_{4n+3}.
$$
Hence, repeating $K$ times,
$$
A_n = \sum_{1 \leq k \leq K} \sum_{0 \leq m < 2^{k-1}}
\frac{1}{(2^k n + 2m)(2^k n + 2m + 1)} 
+ \sum_{0 \leq q < 2^K} A_{2^K n + q}.
$$
Using the bounds $0 < A_n < \frac{1}{n} - \frac{1}{n+1}$ and
telescoping, the last sum is less than $2^{-K}$. Letting $K$
tend to infinity, we obtain the (rapidly convergent) series
$$
A_n = \sum_{k \geq 1} \ \sum_{0 \leq m < 2^{k-1}}
\frac{1}{(2^k n + 2m)(2^k n + 2m + 1)}.
$$
Substituting into the sum defining $S$ yields the double series
$$
S = \sum_{n \geq 1} \ \sum_{k \geq 1} \ \sum_{0 \leq m < 2^{k-1}}
\frac{r_n}{(2^k n + 2m)(2^k n + 2m + 1)},
$$
which converges absolutely. Thus we may collect terms with
the same denominator, and we arrive at the series
$$
S = \sum_{i \geq 1} \frac{R'_i}{2i(2i+1)},
$$
where
$$
R'_{i} := \sum_{n \in {\cal E}_{i}} r_n
$$
with ${\cal E}_{i} := \{n \geq 1, \ \exists k \geq 1, \
\exists m \in [0, 2^{k-1}), \ 2^{k-1}n + m = i \}$.
On the one hand, this proves that the series 
$\sum \frac{R'_i}{2i(2i+1)}$ is absolutely convergent 
(hence the series $\sum |R'_i| i^{-2}$ is convergent).
On the other hand, $R'_i$ can also be written as
$$
R'_{i} := \sum_{1 \leq k \leq \frac{\log i}{\log 2} + 1} 
r_{\lfloor\frac{i}{2^{k-1}}\rfloor} = 
\sum_{k \geq 0} r_{\lfloor\frac{i}{2^k}\rfloor}
$$
(recall that we have set $r_0 := 0$). Thus $R'_i = R_i$ 
by the hypothesis, and the proof is complete. \endpf

\begin{example}\label{fouroverpi}
Theorem~\ref{weight} yields in particular the series for $\gamma$ and 
$\log \frac{4}{\pi}$ in the Introduction, in Example~\ref{gammaplusminus},
and in \cite[Corollary 1]{Sondow2}. Namely,

If $r_1 = r_2 = \cdots = 1$, then the series defining $S$ sums to 
$\gamma$ from Lemma~\ref{mainlemma}, and the formula defining $R'_i=R_{i}$ 
reduces to $R_{i} = \lfloor \frac{\log 2i}{\log 2} \rfloor = 
N_{1,2}(i) + N_{0,2}(i)$.

If $r_n = (-1)^{n-1}$, then $S = \log \frac{4}{\pi}$ 
(see \cite{Sondow1} or decompose $S$ into 
$\sum$ (odd terms) 
$- \sum$ (even terms) and apply Lemma~\ref{mainlemma}), and the 
formula defining $R'_i=R_{i}$ implies $R_{i} = N_{1,2}(i) - N_{0,2}(i)$.
To see this equality, first note that if it holds for $i \geq 1$, then 
using Lemma 5 and looking at the cases $n$ odd and $n$ even,
$$
r_n = R_n - R_{\left\lfloor \frac{n}{2} \right\rfloor} = (-1)^{n-1}
$$
for $n \geq 1$ (compare \cite[Lemma~2]{Sondow2}). Now recall that 
properties (1) and (2) in Lemma~\ref{transform} are equivalent.

\end{example}

\begin{remark}\label{newproof}

Example~\ref{fouroverpi} shows that it is possible to deduce the formula
$$
\sum_{n \geq 1} \frac{N_{1,2}(n)-N_{0,2}(n)}{2n(2n+1)} = \log \frac{4}{\pi}
$$
 from Theorem~\ref{weight} and Lemma~\ref{mainlemma} without using
Lemma~\ref{AS}: this yields a proof of the formula that is different 
from those in \cite{Sondow2} and Example~\ref{gammaplusminus}. Similar 
reasoning applies for any ultimately periodic sequence $(r_n)_{n \geq 1}$. 
In particular, it is not hard to see that the relations giving $r_{2n}$ and 
$r_{2n+1}$ in terms of the $R_i$'s imply that the sequence $(r_n)_{n \geq 1}$ 
is periodic whenever $R_i = N_{w,2}(i)$ for some fixed $w$ and for every 
$i \geq 1$. Hence Theorem~\ref{degree2} can be deduced from Theorem~\ref{weight} 
and Lemma~\ref{mainlemma} (along with the method for decomposing series employed
in Example~\ref{fouroverpi}), without using Lemma~\ref{AS}. In the same vein, 
the generalization in Section~\ref{gal} can be proved using a
generalization of Theorem~\ref{weight} to base $B$ together with
Lemma~\ref{mainlemma}.

\end{remark}

\section{Future directions}

Lemma~\ref{AS} is the main tool for summing series in \cite{AS} and in the 
present paper. It might be possible to use the lemma to obtain the base $B$ 
accelerated series for Euler's constant in \cite[Theorem 2]{Sondow2}, and 
to sum more general series with $N_{w,B}(n)$. On the other hand, it might 
also be possible to extend the results of \cite{AS} and the present paper, 
and sum series where $(N_{w,B}(n))_{n \geq 1}$ is replaced by a more
general integer sequence $(a_n)_{n \geq 1}$, using the decomposition in 
\cite{MM} of a sequence $(a_n)_{n \geq 1}$ into a (possibly infinite) 
linear combination of block-counting sequences $(N_{w,B}(n))_{n \geq 1}$ 
(see also \cite{ASreg}). Of course, since this may replace a series with 
an infinite sum, for the method to work the new series must be summable 
in closed form.

\end{document}